Cassini Significant Events for 03/28/07 - 04/03/07
The most recent spacecraft telemetry was acquired on Tuesday, April 3, from the Goldstone tracking complex. The Cassini spacecraft is in an excellent state of health and all subsystems are operating normally. Information on the present position and speed of the Cassini spacecraft may be found on the “Present Position” web page located at http://saturn.jpl.nasa.gov/operations/present-position.cfm.
Wednesday, March 28 (DOY 087):
The S28 background sequence concluded and S29 began execution today at 2007-087T08:04:00 Spacecraft Event Time. The sequence will run for 37 days and conclude on May 4, 2007. During that time, there will be two targeted encounters with Titan and two non-targeted flybys - one of Dione and one of Telesto. Eight maneuvers are scheduled, numbered 101 through 108. The sequence will be split into two parts with part one uplinked on March 25, and part two uplinked on April 19.
Science early in the sequence included a Composite Infrared Spectrometer (CIRS) compositional map of Titan, and an Imaging Science Subsystem (ISS) photopolarimetry observation of Titan. As the week progressed, the Magnetospheric and Plasma Science (MAPS) instruments performed a magetospheric campaign and took observations of dusk-side magnetospheric boundaries at a variety of latitudes. This was followed by a Visual and Infrared Mapping Spectrometer (VIMS) stellar occultation, Ultraviolet Imaging Spectrograph scans of Enceladus, and an ISS satellite orbit determination observation. The week ended with a VIMS movie of the F ring.
A delivery coordination meeting was held today for Mission Sequence Subsystem (MSS) version D12.1.1. This is a patch delivery incorporating updates to the SSR Management Tool and Sequence Generator to support the swapping of primeness of the SSRs. In addition, updates were made to the Pointing Design Tool, SEG, Sequence Phase List of Ancillary Files, and some run scripts.
Orbit Trim Maneuver (OTM) #101 was performed today. This is the cleanup maneuver for the Titan 27 encounter on March 25. The main engine burn began at 2:59 PM PST. Telemetry immediately after the maneuver showed the burn duration was 3.22 seconds, giving a delta-V of 0.527 m/s. All subsystems reported nominal performance after the maneuver.
Thursday, March 29 (DOY 088):
The Spacecraft Operations Office hosted a review today of all engineering activities to take place during S34. This is the period of time between Sept. 22 and Oct. 31 of 2007.
Science Planning has performed an analysis of Live Inertial Vector Propagator (IVP) Update targeting based on the orbit determination (OD) solution released today from OTM-101. A maximum error of 0.963 mrad for the Dione observation on DOY 98 has been identified. An analysis of the OD accuracy shows Dione to be within the ISS narrow angle camera field of view without an update. The pointing is within the acceptable range for all the active participants - VIMS, CIRS, and ISS. Consequently the S29 Live IVP Update is officially cancelled.
Monday, April 2 (DOY 092):
Orbit Trim Maneuver (OTM) #102 was performed today. This is the apoapsis maneuver setting up for the Titan 28 encounter on April 10. The main engine burn began at 10:45 PM PST. Telemetry immediately after the maneuver showed the burn duration was 16.7 seconds, giving a delta-V of 2.69 m/s. All subsystems reported nominal performance after the maneuver.
The eighth scheduled archive delivery of Cassini science data products occurred on April 1. Most teams delivered to the Planetary Data System early to support the Cassini Data Analysis Program (CDAP). Of note, Radar delivered T16 data, originally scheduled for delivery on July 1, 2007. Upcoming archive deliveries in July and October will include Radar altimetry data, Cosmic Dust Analyzer-High Rate Detector data, and updates to the MAG, Ultraviolet Imaging Spectrograph, and Magnetospheric Imaging Instrument calibration files.
An encounter strategy meeting was held today to cover the period between April 10 and April 26, Titan flybys T28 and T29, and maneuvers 104-106.
Tuesday, April 3 (DOY 093):
The Science Operations Plan update process for S31 completed on Friday, Mar. 30. Products were handed off to the sequence leads and the final development process for S31 kicked off today. The meeting included a discussion of designating SSR-B as the prime SSR for recording data and using SSR-A for high value data recording as needed. The recommendation was approved, but may be revisited prior to the installation of a CDS flight software update in S34. Merged sequence products and stripped subsequences have been published to the file repository for team review.
The hexagonal cloud system on Saturn is Astronomy Picture of the Day today.
Check out the Cassini web site at http://saturn.jpl.nasa.gov for the latest press releases and images.